OpenStudy (anonymous):

Which is another way to write the expression 25 • (r + 6)? A. (25 • r) + (25 • 6) B. (25 • r) + 6 C. 25 • r + 6 D. 25 • r • 6

OpenStudy (anonymous):

Using this, can you tell me the answer? \[a*(b+c) = a*b + a*c\]

OpenStudy (anonymous):

its definitely not D or C so by using the process of elimination

OpenStudy (anonymous):

Well using what I posted and letting a=25, b=r and c=6. What is the answer? \[a*(b+c) = a*b + a*c\]\[25*(r+6) = ?\]